OpenStudy (anonymous):

Consider the line that passes through the points (2, 3) and (6, 11). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y -11 = 2(x - 6) y - 3 = 2(x - 2)

OpenStudy (anonymous):

A.Only equation #1 represents the line that passes through the two given points. B. Only equation #2 represents the line that passes through the two given points. C. Both equation #1 and equation #2 represent the line that passes through the two given points. D. neither equation #1 nor equation #2 represents the line that passes through the two given points

OpenStudy (texaschic101):

eq 1 : y - 11 = 2(x - 6) eq 2 : y - 3 = 2(x - 2) do you know how to find the slope ?

OpenStudy (texaschic101):

slope (m) = (y2 - y1) / (x2 - x1) (2,3) x1 = 2, y1 = 3 (6,11) x2 = 6, y2 = 11 now plug these into your formula and find the slope

OpenStudy (texaschic101):

hello ???

OpenStudy (anonymous):

yess i found out barely thanks

OpenStudy (texaschic101):

did you find the slope ?

OpenStudy (anonymous):

yeah its B right or unless i did it wrong

OpenStudy (texaschic101):

let me check....hold on

OpenStudy (texaschic101):

m= (y2 - y1) / (x2 - x1) m = (11 - 3) / (6 - 2) m = 8/4 slope (m) = 2 now put it in this formula y - y1 = m(x - x1) using points (2,3) and slope 2 y - 3 = 2(x - 2) same thing here...using points(6,11) and slope 2 y - y1 = m(x - x1) y - 11 = 2(x - 6) wrong....answer is not b it is c....they both work

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (anonymous):

nope now i get it thankss :)

OpenStudy (texaschic101):

your welcome :)